A real value function f ( x ) satisfied the functional equation f ( x-y ) = f ( x ) f ( y ) - f ( a-x ) f ( a + y ) and f ( 0 ) = 1, f ( 2a-x ) =? 
 

Grade:12th pass

1 Answers

Riddhish Bhalodia
8 years ago
Put x=y=0 and then we get
f(a) = 0
now put x=a
then
f(a-y) = f(a)f(y) – f(0)f(a+y) = -f(a+y)
now put
y = x-a
then
f(2a-x) = -f(x)